OpenStudy (anonymous):

slope 5/2; passes through (0,2)

OpenStudy (ivettef365):

what is the question ?

OpenStudy (anonymous):

need to graph this slope 5/2 passes through (0,2)

OpenStudy (anonymous):

no book sent home dont understand

OpenStudy (anonymous):

use the formula \[y - y_1 = m(x-x_1)\] where your slope is your m and your x and y coordinates are you x1 and y1 and then solve for y and you will get a y =mx +b format

OpenStudy (anonymous):

not sure that helps as I need to graph this

OpenStudy (anonymous):

if all you do is plott the line and not put it into a y=mx + b then all you have to do is go to the coordinate (0.2) then go up 5 spaces then left 2 places do that 3 times..

OpenStudy (anonymous):

why 3 times?

OpenStudy (anonymous):

well you can do it as much as you like just 3 points of the line makes teacher happy with the line..

OpenStudy (anonymous):

lol thank you

OpenStudy (anonymous):

no problem :)